OpenStudy (anonymous):

There is a large bag with marbles in it. There are 23 red ones, 16 black ones, 5 blue ones, 17 orange ones, 4 white ones, and 15 purple ones. Juanita reaches in without looking and selects one marble. What is the probability that it is purple?

hero (hero):

Hint: Add 23 + 16 + 5 + 17 + 4 + 15 to find the total number of marbles.

OpenStudy (anonymous):

80

hero (hero):

Okay, so since \[\text{Probability} = \frac{\text{Part}}{\text{Whole}} =\frac{\text{total number of purple}}{\text{total Number of marbles}} \] What do you believe is the fraction?

OpenStudy (anonymous):

1/15 @Hero

hero (hero):

So there's 1 purple marble and 15 total marbles in the bag?

OpenStudy (anonymous):

no 80 marbles @Hero

hero (hero):

And how many purple marbles are there?

OpenStudy (anonymous):

15

hero (hero):

Okay so the initial fraction should be \(\frac{15}{80}\) no?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

@Hero noo

hero (hero):

Why not?

OpenStudy (anonymous):

@Hero because its not one of the answers

hero (hero):

That's because you haven't reduced it yet.

OpenStudy (anonymous):

ohhh yeahh how do i do that @Hero

hero (hero):

Break down the numerator and denominator into prime factors, then cancel factors of 1.

OpenStudy (anonymous):

1/80 or 1/15? @Hero

hero (hero):

neither. Nice attempt at guessing though.

hero (hero):

You have no clue how to reduce fractions do you?

OpenStudy (anonymous):

not guessing those are just options from my stuff and no not really lol

hero (hero):

I'm sorry to hear that.

OpenStudy (anonymous):

thanks lol